OpenStudy (anonymous):

if i = sq root of -1 what is the value of i^4

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[ \sqrt{-1}=i\implies i^2=-1 \]

OpenStudy (anonymous):

\[ i^4 = (i^2)^2=(-1)^2 \]

geerky42 (geerky42):

\[i = \sqrt{-1}\]\[i^2 = -1\]\[i^4 = (i^2)^2 = (-1)^2 = \boxed{1}\]

OpenStudy (anonymous):

\[ -1\times -1 = ? \]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

thanks

geerky42 (geerky42):

You're welcome.